loldiv:

cospi/4cospi/6=1/2(---pi/12+cos5pi/12)

Tranquility:

What is the question asking you to do exactly? Do you know these identities

1 attachment
Tranquility:

For your question, it's derived from the cosine addition and subtraction identities \(\cos(a+b) + \cos(a-b) = 2 \cos (a) \cos (b)\)

1 attachment